### v1

@@ Abstract
Aims. We present an extensive catalog of the physical properties of more than a million galaxies within the Dark Energy Spectroscopic Instrument (DESI), one of the largest spectroscopic surveys to date. Spanning over a full variety of target types, including emission line galaxies and luminous red galaxies as well as quasars, our survey encompasses an unprecedented range of spectroscopic redshifts, stretching from 0 to 6.
Methods. The physical properties, such as stellar masses and star formation rates, are derived via the CIGALE spectral energy distribution (SED) fitting code accounting for the contribution coming from active galactic nuclei (AGN). Based on the modeling of the optical-mid-infrared (grz complemented by WISE photometry) SEDs, we study galaxy properties with respect to their location on the main sequence.
Results. We revise the dependence of stellar mass estimates on model choices and availability of the WISE photometry. The WISE information is mandatory to minimize the misclassification of star-forming galaxies as AGN. The lack of WISE bands in SED fits leads to elevated AGN fractions for 68% of star-forming galaxies identified using emission line diagnostic diagram but does not affect significantly their stellar mass nor star formation estimates.
@@
The exploration of galaxies has been a focal point of astronomical research for centuries, revealing an astonishing diversity of galaxy types and physical processes. Large galaxy photometric catalogs mapping an unprecedented number of galaxies and their properties, such as COSMOS (Scoville et al., 2007; Weaver et al., 2022), CANDELS (Grogin et al., 2011; Koekemoer et al., 2011), UltraVISTA (McCracken et al., 2012; Muzzin et al., 2013), and ZFOURGE (Straatman et al., 2016) complemented by spectroscopic sets, such as Sloan Digital Sky Survey (SDSS; York et al., 2000), Galaxy and Mass Assembly (GAMA; Baldry et al., 2018), Deep Extragalactic VIsible Legacy Survey (DEVILS; Davies et al., 2021), and VIMOS Public Extragalactic Redshift Survey (VIPERS; Scodeggio et al., 2018) have allowed us to unlock fundamental galaxy scaling relations and have made significant contributions to our understanding of galaxies and physical processes regulating their formation and evolution.
@@
Template-based techniques relying on the spectral energy distribution (SED; e.g. Conroy 2013 and references therein) fitting of galaxies are generally used to derive physical properties of galaxies from large-scale sky surveys. These methods heavily depend on the physical models of galaxy populations (e.g. Mitchell et al. 2013; Moustakas et al. 2013; Lower et al. 2020; Pacifici et al. 2023) and a statistical method of finding the best fits (e.g. Leja et al. 2018). The optimization of the parameter space and priors in constructing a model library is vital for finely tuning parameters related to stellar populations, dust content, and other key factors. This precision enables a more accurate characterization of galaxies. Building an extensive variety of models is crucial to encompass the full diversity of galaxy properties, but it introduces challenges such as degeneracies, where different parameter combinations yield similar predictions (e.g. Lower et al. 2020). An excessively large model library also poses the risk of overfitting, where models may fit noise or peculiarities in the data rather than capturing the genuine underlying physical properties of galaxies. SED fitting demands a wide wavelength coverage for better tracing the contribution from all stellar types in a galaxy (e.g. Maraston et al., 2006; Pforr et al., 2019) and to break degeneracies between the host galaxy and the active galactic nuclei (AGN) parameters (e.g. Calistro Rivera et al., 2016; Thorne et al., 2022a). Accounting for the AGN contribution is one of the most significant sources of uncertainty in estimated physical properties as AGN are relatively common and their contribution to the mid-infrared (MIR) emission might be significant (Leja et al., 2018). The inferred physical properties also hinge on the accuracy of photometry and redshift of galaxies (e.g. Acquaviva et al., 2015; Iyer & Gawiser, 2017; Paulino-Afonso et al., 2022). Thus, careful consideration of the extensive range of possible model combinations within the parameter space is essential for the robust estimations of physical properties of galaxies (e.g. Pforr et al., 2012; Johnson et al., 2021; Han et al., 2023; Pacifici et al., 2023).
@@
To address the aftermentioned challenges, a collection of panchromatic SED codes has been developed relying on the reduced () techniques such as MAGPHYS (da Cunha et al., 2008), BEAGLE (Chevallard & Charlot, 2016), Prospector (Johnson et al., 2021), BAGPIPES (Carnall et al., 2018), CIGALE (Boquien et al., 2019), ProSpect (Robotham et al., 2020) and the most recent SED fitting codes based on the Markov Chain Monte Carlo (MCMC) approach, such as MCSED (Bowman et al., 2020), piXedfit (Abdurro’uf et al., 2021), Lightning (Doore et al., 2023), PROVABGS (Hahn et al., 2023a), GalaPy (Ronconi et al., 2024), among others. Employing diverse forward-modeling frameworks or templates, along with a range of Bayesian methods such as MCMC sampling or on a model grid, these codes provide a comprehensive approach to accurately estimate the physical properties of galaxies. We refer the reader to Pacifici et al. (2023) and Best et al. (2023) for a review of the performance of different SED fitting tools.
@@
The evolving landscape of astronomy and the state-of-the-art instruments such as the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al. 2016a, 2022), the Prime Focus Spectrograph (PFS; Takada et al. 2014), the Vera C. Rubin Observatory (Ivezić et al. 2019), the James Webb Space Telescope (Gardner et al. 2006), Euclid (Laureijs et al., 2010), and the Nancy Grace Roman Space Telescope (Spergel et al. 2015), demands an even deeper understanding of the physical properties of galaxies, accounting for various target types and spanning a wider redshift range. In particular, the advent of large multi-wavelength surveys triggered using SED fitting methodology to constrain AGN and its host galaxy properties for statistical samples (e.g. Walcher et al., 2011; Boquien et al., 2019; Johnson et al., 2021; Yang et al., 2020, 2022; Thorne et al., 2022b; Bichang’a et al., 2024). The importance of incorporating AGN templates for reliable estimates of physical properties of galaxies hosting AGN was already raised in e.g. Ciesla et al. (2015). The SED fitting approach revealed the potential not only to derive reliable properties of AGN and host properties (e.g. Marshall et al. 2022; Mountrichas et al. 2021a; Burke et al. 2022; Best et al. 2023) but also to identify AGN based on their multi-wavelength information (e.g. Thorne et al. 2022b; Best et al. 2023; Yang et al. 2023; Prathap et al. 2024). AGN SED modeling techniques are also used as the base for the target selection of forthcoming wide-field spectroscopic surveys such as 4MOST (Merloni et al. 2019) and VLT-MOONS (Maiolino et al. 2020).
@@
In this paper, we describe the methodology employed in constructing the Value Added Catalog (VAC) of physical properties of DESI Early Data Release (EDR) galaxies obtained via SED modeling with the Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019). This code, based on the energy balance principle, has already proved its efficiency and accuracy in estimating physical properties accounting for the AGN contribution (e.g. Ciesla et al., 2015; Yang et al., 2023). Its modular framework allows the inclusion of various AGN models, both based on both theoretical approaches (e.g. Fritz et al., 2006) and observational constraints (e.g. Stalevski et al., 2012, 2016). We demonstrate the utility of our catalog by showcasing its potential for discriminating host galaxy properties as well as investigate the influence of the model assumptions and photometry availability. In the follow-up paper (Siudek et al. under DESI Collaboration review), we discuss the ability of the SED fitting approach to distinguish narrow-line (NL) and broad-line (BL) AGN based on their physical properties. The structure of the paper is as follows. In Sect. 2, we provide an overview of the DESI survey and EDR data. In Sect. 3, we describe the SED fitting methodology applied to derive the physical properties of DESI galaxies. The general properties of the catalog are presented in Sect. 4. To validate the derived properties of galaxies we compared to existing catalogs in Sec. 5. In Sec. 6 we discuss the dependence of the physical properties on model assumptions and MIR availability. Finally, Sect. 7 summarizes the catalog and our analysis.
@@ 2 DESI data
DESI is a 5000-fiber multiobject spectrograph installed on the Mayall 4-meter telescope at Kitt Peak National Observatory. It covers a spectral range of Å with a wavelength-dependent spectral resolution, (DESI Collaboration et al., 2016b, 2022; Miller et al., 2023; Silber et al., 2023). It is designed to observe approximately 36 million galaxies (Hahn et al., 2023b; Raichoor et al., 2023; Zhou et al., 2023) and 3 million quasars (Chaussidon et al., 2023) over a 14,000 within a five-year period (DESI Collaboration et al., 2023) with the aim to determine the nature of dark energy through the most precise measurement of the expansion history of the universe ever obtained (Levi et al., 2013). The DESI dataset will be ten times larger than the SDSS (York et al., 2000; Almeida et al., 2023) sample of extragalactic targets and substantially deeper than prior large-area surveys (DESI Collaboration et al., 2023). In December 2020, DESI started a 5-month Survey Validation (SV) before the start of the Main Survey (DESI Collaboration et al., 2024a). The SV campaign consisted of three phases: i) SV1: validating the target selections of the five primary target classes: Milky Way Survey (MWS; Cooper et al. 2023), Bright Galaxy Survey (BGS; Hahn et al. 2023b), Luminous Red Galaxies (LRG; Zhou et al. 2023), Emission Line Galaxies (ELG; Raichoor et al. 2023) and Quasars (QSO; Chaussidon et al. 2023), complemented by secondary fiber targets for additional science goals (SCND; e.g. Darragh-Ford et al. 2023; Fawcett et al. 2023. More details regarding the DESI targeting are described in Myers et al. 2023), ii) SV2: operation developments, and iii) SV3 (One-Percent Survey): further optimizing the observing procedures (Schlafly et al., 2023) with very high fiber assignments completeness over an area of 200 , i.e. of 1% of the final DESI Main Survey. The entire SV data, internally known as Fuji, is publicly released as the DESI Early Data Release (EDR; DESI Collaboration et al., 2023) and is used for generating the VAC of physical properties of DESI galaxies presented in this paper. The First Data Release (DR1, DESI Collaboration et al. in prep.) is planned to be released at the first half of 2025. The DR1 already showcases the DESI potential in science Key Papers presenting the two-point clustering measurements and validation (DESI Collaboration et al. in prep.), BAO measurements from galaxies and quasars (DESI Collaboration et al., 2024d), and from the Lya forest (DESI Collaboration et al., 2024c), as well as a full-shape study of galaxies and quasars (DESI Collaboration et al. in prep). These are complemented by the cosmological results from the BAO measurements (DESI Collaboration et al., 2024b) and the full-shape analysis (DESI Collaboration et al. in prep.), as well as constraints on primordial non-gaussianities (DESI Collaboration et al. in prep.).
@@
DESI spectra are processed with a fully automatic pipeline (Guy et al., 2023), followed by the spectral classification and redshift estimation using the Redrock pipeline111https://github.com/desihub/redrock (Anand et al. 2024, Bailey et al. in prep.). This method relies on the principal component analysis templates generated from a combination of real and synthetic spectra of astronomical sources using an iterative principal component generator (Bailey, 2012), which also takes uncertainties of the data into account. Along with the redshift (Z), redshift uncertainty (ZERR), a redshift warning bitmask (ZWARN), Redrock also assigns a spectral type (SPECTYPE) to every target based on the best fit. The resulting DESI EDR redshift catalog consists of 2,847,435 spectra of 2,757,937 unique sources (DESI Collaboration et al., 2023). For multiply-observed targets, we choose the ’best spectrum’ as the one which has a higher signal to noise ratio (SNR) spectrum, along with good fiber and redshift measurements (ZCAT_PRIMARY = True222https://github.com/desihub/desispec/blob/0.51.13/py/desispec/zcatalog.py#L13). Furthermore, we select sources that are assigned as GALAXY or QSO by Redrock that do not have any fiber issues (COADD_FIBERSTATUS = 0333https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#spectroscopic-reduction-masks) and have a reliable redshift (ZWARN = 0 or 4444https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#redshift-fitting-redrock-masks). We refer the reader to DESI Collaboration et al. 2023 for more details about selection choices. After applying all these cuts, we have a sample of 1,345,137 objects spanning a redshift range of over . Our analysis requires photometric measurements, which are described in the following Subsection.
@@ 2.1 Photometry
DESI primary targets (MWS, BGS, LRG, ELG, and QSO) are selected from the ninth data release of the Legacy imaging surveys555https://www.legacysurvey.org/dr9/description/ (LS/DR9; Dey et al., 2019), which is a combination of three public projects. The northern sky is covered by the Beijing-Arizona Sky Survey (BASS; Zou et al. 2017) in the g and r band and by the Mayall z-band Legacy Survey (MzLS) in the z band with a 5 detection limit of g = 23.48, r = 22.87, and z = 22.29 AB magnitude. The south LS footprint is mapped by the Dark Energy Camera Legacy Survey (DECaLS) in all three bands (g, r, and z) with a 5 detection of g = 23.72, r = 23.27, and z = 22.22 AB magnitude. The detection limits are found for a fiducial galaxy size of .
The ground-based optical and near-infrared (NIR) photometry (i.e., grz photometry, which we will hereafter shortly refer to as optical) is complemented by observations from MIR bands at and 22 m provided by the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010) and a mission extension NEOWISE-Reactivation forced-photometry (Mainzer et al., 2014) in the unWISE maps at the DESI footprint (Meisner et al., 2017; Schlafly et al., 2019; Meisner et al., 2021a). The WISE photometry is matched to optical imaging using the Tractor package (Lang et al., 2016) based on seven-year custom stacks of WISE/NEOWISE exposures, called unWISE coadds, reaching limiting magnitudes of 21.7 and 20.9 AB mag in W1 and W2 (Meisner et al., 2019, 2021b). Table 1 summarizes the information about photometric bands used in this analysis.
@@
| W1 | unWISE | 16.7 |
| W2 | unWISE | 14.5 |
Tractor is a pioneered Python tool based on the statistically rigorous forward-modeling approach to perform source extraction on the pixel data. It is designed to fit images and photometry to estimate source shapes and brightness properties taking into account their different point spread function (PSF) and different band sensitivities. This approach is particularly useful to process LS sources given their wide range of PSF shapes and sizes: the optical data have a typical PSF of arcsec; and the WISE PSF full width at half maximum (FWHM) is arcsec in W1–W3 and arcsec in W4 (Dey et al., 2019). The pixels associated with each detection (called blob) are fitted with models of surface brightness, including the Sérsic profile, and the best fit is chosen as the one which minimizes the . The fits are performed separately on each photometric band (g, r, and z bands) accounting for different PSF and sensitivity of each image. In DR9 the light profiles are fitted with four models: point source (PSF), round exponential galaxy model (REX), de Vaucouleurs model (DEV), exponential model (EXP), and a Sérsic model (SER)666Note that the Sérsic model replaced the composite model used in the previous, DR8, release., in that order. However, Tractor models do not include more complex structures, and the resulting models may not always be ideal. The best-fit model is determined by convolving each model with the specific PSF for each exposure, fitting it to each image, and minimizing the residuals for all images.
@@
The Tractor model fits are determined using only the optical (grz) data. The MIR photometry for each optically detected source is then determined by forcing the location and shape of the model, convolving with the WISE PSF, and fitting to the WISE stacked image. The advantage of the ’forced photometry’ is the ability to deblend any confused WISE sources by using the higher-spatial-resolution optical data and detecting fainter sources than traditional approaches while preserving the photometric reliability. However, this approach limits the LS catalog to only contain WISE photometry for sources that are detected at optical wavelengths (Dey et al., 2019).
Tractor returns the object position, fluxes, galactic extinctions, shape, and morphological parameters, given by the Sérsic index among others. The Tractor catalog also contains a set of quality measures, like FRACMASKED, FRACFLUX, and FRACIN that quantify the quality of the data in each profile fit and can be used to preselect high-quality samples (see Sec. 4.1). More detail about the data reduction can be found in Dey et al. (2019) and Schlegel et al. in prep.
In addition to the primary targets, DESI EDR also includes some secondary targets and targets of opportunity that do not have LS/DR9 photometry. Given that our analysis requires photometric measurements, we only consider the DESI sources that have LS/DR9 photometry (RELEASE = 9010 or 9011 or 9012). This leads to a final sample of 1,286,124 unique objects spanning a redshift range of . The demographics of the the DESI primary targets (MWS, BGS, LRG, ELG, and QSO) is summarized in Tab. 2. The redshift distribution of all these sources, along with their distributions from different targeting types is shown in Fig. 1 with the distribution of MWS scaled up by a factor of ten to more easily compare the shape of each distribution. The DESI VAC covers a wide redshift range spanning over targeting BGS galaxies and AGN at lower redshift (; Hahn et al. 2023b; Juneau et al. 2024) and ELGs at higher redshift (; Raichoor et al. 2023) with QSO spanning out to (Chaussidon et al. 2023). LRGs extend out to (Zhou et al. 2023) and SCND sources cover the entire redshift range incorporating low-z targets (e.g. Darragh-Ford et al., 2023) as well as high-z QSOs (e.g. Fawcett et al., 2023). The catalog also includes a negligible fraction ( of galaxies and QSO observed within the MWS (Cooper et al. 2023) spanning over a wide redshift range. Namely, it is composed of 3,238 galaxies (with a mean redshift ) and 6,052 QSO (with a mean redshift ).

### v2

@@ Abstract
Aims. We present an extensive catalog of the physical properties of more than a million galaxies investigated with the Dark Energy Spectroscopic Instrument (DESI), one of the largest spectroscopic surveys to date. Spanning a full range of target types, including emission-line galaxies, luminous red galaxies, and quasars, our survey encompasses an unprecedented range of spectroscopic redshifts, all the way from 0 to 6.
Methods. The physical properties, such as stellar masses and star formation rates, were derived via the CIGALE spectral energy distribution (SED) fitting code accounting for the contribution coming from active galactic nuclei (AGNs). Based on the modeling of the optical-mid-infrared (grz supplemented with WISE photometry) SEDs, we studied the galaxy properties with respect to their location on the main sequence.
Results. We have revised the dependence of stellar mass estimates on model choices and on the availability of WISE photometry. Indeed, the WISE data are required to minimize the misclassification of star-forming galaxies as AGNs. The lack of WISE bands in SED fits leads to elevated AGN fractions for 68% of star-forming galaxies identified using emission line diagnostic diagrams, but this does not significantly affect their stellar mass or star formation estimates.
@@
The exploration of galaxies has been a focal point of astronomical research for centuries, revealing an astonishing diversity of galaxy types and physical processes. Large galaxy photometric catalogs mapping an unprecedented number of galaxies and their properties, such as COSMOS (Scoville et al., 2007; Weaver et al., 2022), CANDELS (Grogin et al., 2011; Koekemoer et al., 2011), UltraVISTA (McCracken et al., 2012; Muzzin et al., 2013), and ZFOURGE (Straatman et al., 2016) supplemented with spectroscopic datasets, such as those of Sloan Digital Sky Survey (SDSS; York et al., 2000), Galaxy and Mass Assembly (GAMA; Baldry et al., 2018), Deep Extragalactic VIsible Legacy Survey (DEVILS; Davies et al., 2021), and VIMOS Public Extragalactic Redshift Survey (VIPERS; Scodeggio et al., 2018) have allowed us to unlock fundamental galaxy scaling relations and have made significant contributions to our understanding of galaxies and physical processes regulating their formation and evolution.
@@
Template-based techniques relying on the spectral energy distribution (SED; e.g., Conroy 2013 and references therein) fitting of galaxies are generally used to derive physical properties of galaxies from large-scale sky surveys. These methods heavily depend on the physical models of galaxy populations (e.g., Mitchell et al. 2013; Moustakas et al. 2013; Lower et al. 2020; Pacifici et al. 2023) and a statistical method of finding the best fits (e.g., Leja et al. 2018). The optimization of the parameter space and priors in constructing a model library is vital for finely tuning parameters related to stellar populations, dust content, and other key factors. This precision enables a more accurate characterization of galaxies. Building an extensive variety of models is crucial to encompass the full diversity of galaxy properties, but it introduces challenges such as degeneracies, where different parameter combinations yield similar predictions (e.g., Lower et al. 2020). An excessively large model library also poses the risk of overfitting, whereby models may fit noise or peculiarities in the data, rather than capturing the genuine underlying physical properties of galaxies. The SED fitting demands a wide wavelength coverage for better tracing the contribution from all stellar types in a galaxy (e.g., Maraston et al., 2006; Pforr et al., 2019) and to break degeneracies between the host galaxy and the parameters related to the active galactic nucleus (AGN) (e.g., Calistro Rivera et al., 2016; Thorne et al., 2022a). Accounting for the AGN contribution is one of the most significant sources of uncertainty in estimated physical properties as AGNs are relatively common and their contribution to the mid-infrared (MIR) emission might be significant (Leja et al., 2018). The inferred physical properties also hinge on the accuracy of photometry and redshift of galaxies (e.g., Acquaviva et al., 2015; Iyer & Gawiser, 2017; Paulino-Afonso et al., 2022). Thus, a careful consideration of the extensive range of possible model combinations within the parameter space is essential for the robust estimations of physical properties of galaxies (e.g., Pforr et al., 2012; Johnson et al., 2021; Han et al., 2023; Pacifici et al., 2023).
@@
To address the aftermentioned challenges, a collection of panchromatic SED codes has been developed relying on the reduced () techniques such as MAGPHYS (da Cunha et al., 2008), BEAGLE (Chevallard & Charlot, 2016), Prospector (Johnson et al., 2021), BAGPIPES (Carnall et al., 2018), CIGALE (Boquien et al., 2019), and ProSpect (Robotham et al., 2020), along with the most recent SED fitting codes based on the Markov chain Monte Carlo (MCMC) approach, such as MCSED (Bowman et al., 2020), piXedfit (Abdurro’uf et al., 2021), Lightning (Doore et al., 2023), PROVABGS (Hahn et al., 2023a), and GalaPy (Ronconi et al., 2024), among others. Employing diverse forward-modeling frameworks or templates, along with a range of Bayesian methods such as MCMC sampling or on a model grid, these codes provide a comprehensive approach to accurately estimate the physical properties of galaxies. We refer to Pacifici et al. (2023) and Best et al. (2023) for a review of the performance of different SED fitting tools.
@@
The evolving landscape of astronomy and the state-of-the-art instruments, including Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al. 2016a, 2022), Prime Focus Spectrograph (PFS; Takada et al. 2014), Vera C. Rubin Observatory (Ivezić et al. 2019), James Webb Space Telescope (Gardner et al. 2006), Euclid (Laureijs et al., 2010), and Nancy Grace Roman Space Telescope (Spergel et al. 2015), demands an even deeper understanding of the physical properties of galaxies. We must also account for various target types, spanning a wider redshift range. In particular, the advent of large multi-wavelength surveys triggered using SED fitting methodology to constrain the AGN and its host galaxy properties for statistical samples (e.g., Walcher et al., 2011; Boquien et al., 2019; Johnson et al., 2021; Yang et al., 2020, 2022; Thorne et al., 2022b; Bichang’a et al., 2024). The importance of incorporating AGN templates for reliable estimates of physical properties of galaxies hosting AGNs was already raised in previous works, such as Ciesla et al. (2015). The SED fitting approach has revealed the potential to not only derive reliable properties of AGN and host properties (e.g., Marshall et al. 2022; Mountrichas et al. 2021a; Burke et al. 2022; Best et al. 2023) but also to identify AGNs based on their multi-wavelength information (e.g., Thorne et al. 2022b; Best et al. 2023; Yang et al. 2023; Prathap et al. 2024). The AGN SED modeling techniques are also used as the base for the target selection of forthcoming wide-field spectroscopic surveys such as 4MOST (Merloni et al. 2019) and VLT-MOONS (Maiolino et al. 2020).
@@
In this paper, we describe the methodology employed in constructing the Value Added Catalog (VAC) of physical properties of DESI Early Data Release (EDR) galaxies obtained via SED modeling with the Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019). This code, based on the energy balance principle, has already proved its efficiency and accuracy in estimating physical properties accounting for the AGN contribution (e.g., Ciesla et al., 2015; Yang et al., 2023). Its modular framework allows the inclusion of various AGN models, both based on both theoretical approaches (e.g., Fritz et al., 2006) and observational constraints (e.g., Stalevski et al., 2012, 2016). We demonstrate the utility of our catalog by showcasing its potential for discriminating host galaxy properties, while also investigating the influence of the model assumptions and availability of photometry data. In the follow-up paper (Siudek et al. under DESI Collaboration review), we discuss the ability of the SED fitting approach to distinguish narrowline (NL) and broadline (BL) AGNs based on their physical properties. The structure of the paper is as follows. In Sect. 2, we provide an overview of the DESI survey and EDR data. In Sect. 3, we describe the SED fitting methodology applied to derive the physical properties of DESI galaxies. The general properties of the catalog are presented in Sect. 4. We compared our sample to existing catalogs to validate the derived properties of galaxies in Sect. 5. In Sect. 6, we discuss the dependence of the physical properties on model assumptions and MIR availability. Finally, Sect. 7 summarizes the catalog and our analysis. Throughout this paper, we assume WMAP7 cosmology (Komatsu et al., 2011), with = 0.272 and = 70.4. We also consider the photometry in AB magnitudes (Oke & Gunn 1983).
@@ 2.1 DESI survey
DESI is a 5000-fiber multiobject spectrograph installed on the Mayall 4-meter telescope at Kitt Peak National Observatory. It covers a spectral range of Å with a wavelength-dependent spectral resolution, (DESI Collaboration et al., 2016b, 2022; Miller et al., 2024; Silber et al., 2023). It is designed to observe approximately 36 million galaxies (Hahn et al., 2023b; Raichoor et al., 2023; Zhou et al., 2023) and 3 million quasars (Chaussidon et al., 2023) over a 14,000 within a five-year period (DESI Collaboration et al., 2024a) with the aim to determine the nature of dark energy through the most precise measurement of the expansion history of the universe ever obtained (Levi et al., 2013). The DESI dataset will be ten times larger than the SDSS (York et al., 2000; Almeida et al., 2023) sample of extragalactic targets and substantially deeper than prior large-area surveys (DESI Collaboration et al., 2024a). In December 2020, DESI started a five-month survey validation (SV) before the start of the main survey (DESI Collaboration et al., 2024b). The SV campaign consisted of three phases: i) SV1: validating the target selections of the five primary target classes: Milky Way survey (MWS; Cooper et al. 2023) and bright galaxy survey (BGS; Hahn et al. 2023b), along with surveys of luminous red galaxies (LRG; Zhou et al. 2023), emission-line galaxies (ELG; Raichoor et al. 2023), and quasars (QSO; Chaussidon et al. 2023). These targets were further supplemented with secondary fiber targets for additional science goals (SCND; e.g., Darragh-Ford et al. 2023; Fawcett et al. 2023. More details regarding the DESI targeting are described in Myers et al. 2023); Furthermore, ii) SV2: operation developments; and iii) SV3 (One-Percent survey) optimized the observing procedures (Schlafly et al., 2023), with very high fiber assignments completeness over an area of 200 , namely, of 1% of the final DESI main survey. The entire SV data, internally known as Fuji, is publicly released as the DESI Early Data Release (EDR; DESI Collaboration et al., 2024a) and is used for generating the VAC of physical properties of DESI galaxies presented in this paper. The First Data Release (DR1, DESI Collaboration et al. in prep.) is planned to be released at the first half of 2025. The DR1 already showcases the DESI potential in science Key Papers presenting the two-point clustering measurements and validation (DESI Collaboration et al. in prep.), BAO measurements from galaxies and quasars (DESI Collaboration et al., 2024e), and from the Lya forest (DESI Collaboration et al., 2024d), as well as a full-shape studies of galaxies and quasars (DESI Collaboration et al. in prep). These are supplemented with the cosmological results from the BAO measurements (DESI Collaboration et al., 2024c) and the full-shape analysis (DESI Collaboration et al. in prep.), as well as constraints on primordial non-gaussianities (DESI Collaboration et al. in prep.).
@@
The DESI spectra are processed with a fully automatic pipeline (Guy et al., 2023), followed by the spectral classification and redshift estimation using the Redrock pipeline111https://github.com/desihub/redrock (Anand et al. 2024, Bailey et al. in prep.). This method relies on the principal component analysis templates generated from a combination of real and synthetic spectra of astronomical sources using an iterative principal component generator (Bailey, 2012), which also takes the uncertainties of the data into account. Along with the redshift (Z), we take the redshift uncertainty (ZERR), a redshift warning bitmask (ZWARN), Redrock also assigns a spectral type (SPECTYPE) to every target based on the best fit. The resulting DESI EDR redshift catalog consists of 2,847,435 spectra of 2,757,937 unique sources (DESI Collaboration et al., 2024a). For multiply observed targets, we chose the ”best” spectrum as the one that has a higher signal-to-noise ratio (S/N) spectrum, along with good fiber and redshift measurements (ZCAT_PRIMARY = True222https://github.com/desihub/desispec/blob/0.51.13/py/desispec/zcatalog.py#L13). Furthermore, we selected sources that have been assigned as GALAXY or QSO by Redrock and that do not have any fiber issues (COADD_FIBERSTATUS = 0333https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#spectroscopic-reduction-masks) , but do have a reliable redshift (ZWARN = 0 or 4444https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#redshift-fitting-redrock-masks). We refer to DESI Collaboration et al. 2024a for more details about our selection choices. After applying all these cuts, we were left with a sample of 1,345,137 objects spanning a redshift range of over . Our analysis requires photometric measurements, which are described in the following subsection.
@@ 2.2 Photometry
The DESI primary targets (MWS, BGS, LRG, ELG, and QSO) were selected from the ninth data release of the Legacy imaging surveys555https://www.legacysurvey.org/dr9/description/ (LS/DR9; Dey et al., 2019), which is a combination of three public projects. The northern sky is covered by the Beijing-Arizona Sky Survey (BASS; Zou et al. 2017) in the g and r band and by the Mayall z-band Legacy Survey (MzLS) in the z band with a 5 detection limit of g = 23.48, r = 22.87, and z = 22.29 AB magnitude. The south LS footprint is mapped by the Dark Energy Camera Legacy Survey (DECaLS) in all three bands (g, r, and z) with a 5 detection of g = 23.72, r = 23.27, and z = 22.22 AB magnitude. The detection limits are found for a fiducial galaxy size of .
The ground-based optical and near-infrared (NIR) photometry (i.e., grz photometry, which we hereafter shortly refer to as optical) is supplemented with observations from MIR bands at and 22 m provided by the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010) and a mission extension NEOWISE-Reactivation forced-photometry (Mainzer et al., 2014) in the unWISE maps at the DESI footprint (Meisner et al., 2017; Schlafly et al., 2019; Meisner et al., 2021a). The WISE photometry is matched to optical imaging using the Tractor package (Lang et al., 2016) based on seven-year custom stacks of WISE/NEOWISE exposures, called unWISE coadds, reaching limiting magnitudes of 21.7 and 20.9 AB mag in W1 and W2 (Meisner et al., 2019, 2021b). Table 1 summarizes the information about photometric bands used in this analysis.
@@
| W3 | unWISE | 16.7 |
| W4 | unWISE | 14.5 |
Tractor is a pioneered Python tool based on the statistically rigorous forward-modeling approach to perform source extraction on the pixel data. It is designed to fit images and photometry to estimate source shapes and brightness properties taking into account their different point spread function (PSF) and different band sensitivities. This approach is particularly useful to process LS sources given their wide range of PSF shapes and sizes: the optical data have a typical PSF of arcsec; and the WISE PSF full width at half maximum (FWHM) is arcsec in W1–W3 and arcsec in W4 (Dey et al., 2019). The pixels associated with each detection (called blob) are fitted with models of surface brightness, including the Sérsic profile, and the best fit is chosen as the one which minimizes the . The fits are performed separately on each photometric band (g, r, and z bands) accounting for different PSF and sensitivity of each image. In DR9, the light profiles are fitted with four models: point source (PSF), round exponential galaxy model (REX), de Vaucouleurs model (DEV), exponential model (EXP), and a Sérsic model (SER)777Note: the Sérsic model replaced the composite model used in the previous, DR8, release., in that order. However, Tractor models do not include more complex structures and the resulting models may not always be ideal. The best-fit model is determined by convolving each model with the specific PSF for each exposure, fitting it to each image, and minimizing the residuals for all images.
@@
The Tractor model fits are determined using only the optical (grz) data. The MIR photometry for each optically detected source is then determined by forcing the location and shape of the model, convolving with the WISE PSF, and fitting to the WISE stacked image. The advantage of the ”forced photometry” is the ability to deblend any confused WISE sources by using the higher-spatial-resolution optical data and detecting fainter sources than traditional approaches, while preserving the photometric reliability. However, this approach limits the LS catalog to strictly WISE photometry for sources detected at optical wavelengths (Dey et al., 2019).
Tractor returns the object position, fluxes, galactic extinctions, shape, and morphological parameters, given by the Sérsic index (among others). The Tractor catalog also contains a set of quality measures, such as FRACMASKED, FRACFLUX, and FRACIN, which offer a qualitative calculation of the data in each profile fit and can be used to preselect high-quality samples (see Sect. 4.1). More detail about the data reduction can be found in Dey et al. (2019) and Schlegel et al. in prep.
In addition to the primary targets, DESI EDR also includes some secondary targets and targets of opportunity that do not have LS/DR9 photometry. Given that our analysis requires photometric measurements, we only consider the DESI sources that have LS/DR9 photometry (RELEASE = 9010 or 9011 or 9012). This leads to a final sample of 1,286,124 unique objects, spanning a redshift range of . The demographics of the the DESI primary targets (MWS, BGS, LRG, ELG, and QSO) is summarized in Table 2. The redshift distribution of all these sources, along with their distributions from different targeting types is shown in Fig. 1, with the distribution of MWS scaled up by a factor of ten to more easily compare the shape of each distribution. The DESI VAC covers a wide redshift range spanning over targeting BGS galaxies and AGNs at lower redshift (; Hahn et al. 2023b; Juneau et al. 2024) and ELGs at higher redshift (; Raichoor et al. 2023), with QSOs spanning out to (Chaussidon et al. 2023); LRGs extend out to (Zhou et al. 2023) and SCND sources cover the entire redshift range, incorporating low-z targets (e.g., Darragh-Ford et al., 2023) as well as high-z QSOs (e.g., Fawcett et al., 2023). The catalog also includes a negligible fraction ( of galaxies and QSO observed within the MWS (Cooper et al. 2023) spanning a wide redshift range; namely, it is composed of 3,238 galaxies (with a mean redshift ) and 6,052 QSOs (with a mean redshift ).